TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x01x15x18x38

RAT Meeting in Point 8 - Jan 15th

PDF RAT_20070115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x15x18x38/RAT_20070115.pdf

MircoCoccoli - 15 Jan 2007 - 18:40

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Topic attachments
I Attachment History Action Size Date Who Comment
PDFpdf RAT_20070115.pdf r1 manage 66.9 K 2007-01-15 - 18:42 MircoCoccoli  
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-01-15 - MircoCoccoli
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login